Runner A crosses the starting line of a marathon and runs at an average pace of 5.6 miles per hour. Half an hour later, Runner B crosses the starting line and runs at an average rate of 6.4 miles per hour. If the length of the marathon is 26.2 miles, which runner will finish ahead of the other?Explain.
Answer by nerdybill(5405)

You will need to apply the distance formula:
d = rt
where
d is distance
r is rate or speed
t is time
.
Rearranging the formula to solve for t:
t = d/r
.
Runner A:
Time it takes for him to finish:
t = 26.2/5.6 = 4.679 hours (to complete race)
.
Runner B:
Time it takes for him to finish:
t = 26.2/6.4 = 4.094 hours
BUT wait, we have to account for the extra .5 hours -- the time it took for him to cross the start line.
t = 4.094 + .5 = 4.594 hours
.
Since Runner B's time is less than Runner A's, we must conclude Runner B will finish first.